A 1.45 mol sample of chromium contains ..... atoms. 1.68 x 10-22 1.25 1022 8.73 x 1023 2.16 x 1025 4.54 x 1025
Answer #1

No. of atoms = moles * Avogadro's number

Avogadro's number = 6.022*10^23

No. of atoms = 1.45*6.022*10^23

= 8.73*10^23 atoms

If you have any query please comment

If you satisfied with the solution please rate it thanks